# On sharpened singular Adams’ type inequalities and applications to $(p,\frac{n}{2})$-biharmonic equations

### v2

@@ 1. Introduction and main results
for all $m\in\mathbb{N}$ with $m<n$ and $\alpha\in[0,n)$. Consequently, if $\beta>\beta_{\alpha,n,m}$, then the supremum (1.10) is no longer finite. Note that in the second order Sobolev space $W^{2,\frac{n}{2}}(\mathbb{R}^{n})$ with $n\geq 4$, the sharp singular Adams’ inequality was investigated in [33] under the following constraint
$\big{\{}u\in W^{2,\frac{n}{2}}(\mathbb{R}^{n}):~{}\|\Delta u\|^{\frac{n}{2}}_{\frac{n}{2}}+\tau\|u\|^{\frac{n}{2}}_{\frac{n}{2}}\leq 1~{}\text{for all}~{}\tau>0\big{\}}.$
@@
Later, this result was again generalized in [52, Theorem 1.1] into the arbitrary Sobolev space $W^{m,\frac{n}{m}}(\mathbb{R}^{n})$. Explicitly, they proved a strengthened version of the singular Adams’ inequality as follows: let $\tau>0$, then
@@
$\sup_{u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}),~{}\|\nabla^{m}u\|^{\frac{n}{m}}_{\frac{n}{m}}+\tau\|u\|^{\frac{n}{m}}_{\frac{n}{m}}~{}\leq 1}~{}\int_{\mathbb{R}^{n}}\frac{\Phi(\beta|u|^{\frac{n}{n-m}})}{|x|^{\alpha}}\,\mathrm{d}x<+\infty,~{}\forall~{}\beta\leq\beta_{\alpha,n,m},$ (1.11)
@@
In recent years, Adams’ type inequalities have been widely studied by many authors across diverse domains such as Hyperbolic spaces, Lorentz spaces, CR spheres, compact Riemannian manifolds, and so on. In this context, we cite some delightful works captured in [28, 27, 26]. For the applications of Adams’ inequalities in the analysis of elliptic PDEs, we also refer to some notable articles [24, 15, 8].
It is worth mentioning that the concentration-compactness principle was first studied in higher-order Sobolev spaces in [23, Theorem 1]. Later, on exploiting the Hilbert structure of $W^{2,2}(\mathbb{R}^{4})$, in [15] authors established a sharp singular concentration-compactness principle and as an application they studied a class of biharmonic equations in dimension four. Then, it was fully extended into whole space $\mathbb{R}^{n}$ in [41]. His result can be read as follows: let $u_{k}\rightharpoonup u$ weakly in $W^{m,\frac{n}{m}}(\mathbb{R}^{n})$ with $u\neq 0$, $\|u_{k}\|_{m,\frac{n}{m}}\leq 1$ and
@@
In addition, he also proved that the above result holds whenever $\|\cdot\|_{m,\frac{n}{m}}$ is replaced by the Ruf norm on $W^{m,\frac{n}{m}}(\mathbb{R}^{n})$, which is defined by
$\|u\|_{\text{Ruf}}:=\big{(}\|u\|^{\frac{n}{m}}_{\frac{n}{m}}+\|\nabla^{m}u\|^{\frac{n}{m}}_{\frac{n}{m}}\big{)}^{\frac{m}{n}},~{}\forall~{}u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}).$
Moreover, we also mention that the asymptotic behavior of the supremum for the subcritical sharp Moser-Trudinger inequalities in $\mathbb{R}^{n}$ with different function spaces can be found in the recent works [36, 35, 45]. In addition, it was also proved in these articles that the critical and subcritical sharp Moser-Trudinger inequalities are indeed equivalent. Similar results for higher order Sobolev spaces in $\mathbb{R}^{n}$ can be seen in [34] and references therein.
To study ($\boldsymbol{S}$), we first define some suitable function spaces. Notice that due to the appearance of two operators in ($\boldsymbol{S}$), we work with the function space
$E:=\Big{\{}u\in L^{1}_{\textit{loc}}(\mathbb{R}^{n})\colon~{}\int_{\mathbb{R}^{n}}|\Delta u|^{p}\,\mathrm{d}x<+\infty,~{}\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}}\,\mathrm{d}x<+\infty\Big{\}},$
where $1<p<\frac{n}{2}$ with $n\geq 4$, which is the completion of the $C_{0}^{\infty}(\mathbb{R}^{n})$ with respect to the following norm
@@
[…]
$\|u\|\leq\|u\|_{E}\leq 2^{\frac{n-2}{n}}\|u\|,~{}\forall~{}u\in E.$ (1.12)
Using standard arguments, we remark that $(E,\|\cdot\|_{E})$ is a uniform convex, reflexive and separable Banach space. Moreover, since reflexivity is preserved under equivalent norms, we infer that $(E,\|\cdot\|)$ is also a reflexive Banach space. In addition, the space $D^{2,p}(\mathbb{R}^{n})$ is defined as the completion of $C_{0}^{\infty}(\mathbb{R}^{n})$ equipped with the norm
@@
Due to the Sobolev embedding, one can see that the chain of embeddings $E\hookrightarrow D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$ hold.
For any $\gamma\in[0,n)$, we define the space $L^{r}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$, consisting of all real-valued measurable functions, with $|u|^{r}|x|^{-\gamma}\in L^{1}(\mathbb{R}^{n})$, equipped with the norm
@@
[…]
  • $o_{k}(1)$ denotes the real sequence converges to zero as $k\to\infty$,
@@
We would like to emphasize the important aspects and novelties of this article below:
@@
    In higher order Sobolev spaces, the Pólya-Szegő types inequalities don’t hold, which causes us to face several barriers in proving the singular Adams’ type inequality in $E$ and the associated concentration-compactness principle, as can be shown in Theorem 1.3 and Theorem 1.4.
@@
    Due to the critical exponential growth of the nonlinearity, there is a healthy competition between the nonhomogeneous $(p,\frac{n}{2})$-biharmonic operator and the nonlinear term because the embedding $E\hookrightarrow L_{\Phi_{\alpha,j_{0}}}(\mathbb{R}^{n})$ is not compact, where $\Phi_{\alpha,j_{0}}(\cdot)$ is the Young function defined in (1.3). Therefore, proving the compactness of Palais-Smale sequences for the variational energy associated with our main problem becomes challenging.
@@
    The study of existence of solutions to our main problem relies on the compact embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ for all $\varrho\geq p^{\ast}$ and $\gamma\in(0,n)$, thanks to Theorem 1.7. It follows that our method is different from $\gamma=0$.
@@
    It is well known that for any $u\in E$, there is no guarantee that $|u|,~{}u^{\pm}\in E$, where $u^{\pm}=\max\{\pm u,0\}$ and thus we shall unable to get positive solutions to our main problem.
@@
Motivated by the above-mentioned works and primarily inspired by the works in [48, 33, 41, 15, 13, 5, 34], we establish a new version of sharp singular Adams’ inequality in the space $E$, its concentration-compactness principle, and as a byproduct of these inequalities, we study a class of elliptic pde involving $(p,\frac{n}{2})$-biharmonic operator in the entire space $\mathbb{R}^{n}$. The following defines the weak formulation of our main problem.
@@
We say that $u\in E$ is a weak solution to the problem ($\boldsymbol{S}$), if there holds
$\int_{\mathbb{R}^{n}}|\Delta u|^{p-2}\Delta u~{}\Delta v~{}\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}-2}\Delta u~{}\Delta v~{}\mathrm{d}x=\int_{\mathbb{R}^{n}}\frac{g(x,u)}{|x|^{\gamma}}v\,\mathrm{d}x,~{}\forall~{}v\in E.$
Briefly speaking, as our interest is to study the problem ($\boldsymbol{S}$) in the space $E$, it should pointed out that, unfortunately, we cannot use the singular Adams’ inequality that appeared in [33, Theorem 1.5]. Hence, we need to demonstrate a version of such inequality.
Theorem 1.3 (Sharp Singular Adams’ Type Inequality).
Let $1<p<\frac{n}{2}$, $0\leq\gamma<n$ and $n\geq 4$. Then for all $0\leq\alpha\leq\beta_{\gamma,n}:=\big{(}1-\frac{\gamma}{n}\big{)}\beta(n,2)$ and $u\in E$, there holds
$\sup_{u\in E,~{}\|u\|\leq 1}\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x<+\infty,$ (1.13)
where $\Phi_{\alpha,j_{0}}(\cdot)$ is defined in (1.3) and $\beta(n,2)=n[(n-2)\omega^{\frac{2}{n}}_{n-1}]^{\frac{n}{n-2}}$ with $\omega_{n-1}$ stands for the measure of the unit sphere in $\mathbb{R}^{n}$. Further, the constant $\beta_{\gamma,n}$ is sharp, that is, if $\alpha>\beta_{\gamma,n}$, then the supremum (1.13) is infinite.
Theorem 1.4 (Concentration-Compactness Principle).
Suppose that $1<p<\frac{n}{2}$, $0\leq\gamma<n$ and $n\geq 4$. Let $\{u_{k}\}_{k}\subset E$ be a sequence such that $\|u_{k}\|=1$ and $u_{k}\rightharpoonup u\neq 0$ in $E$ as $k\to\infty$, then for all $\ell$ satisfying
we have
$\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$ (1.14)
The following natural question is still open at this time:
[…]
Theorem 1.5 ( Subcritical Sharp Singular Adams’ Type Inequality).
Let $1<p<\frac{n}{2}$, $0\leq\gamma<n$, $n\geq 4$ and $0\leq\ell<\beta(n,2)$. Then there holds
The constant $\beta(n,2)$ is sharp in the sense that $\text{ATSC}(\beta(n,2),\gamma)=+\infty$. Also, there exists two positive constants $c(n,\gamma,p)$ and $C(n,\gamma,p)$ such that whenever $l\underset{\approx}{<}\beta(n,2)$,
$\cfrac{c(n,\gamma,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}}\leq\text{ATSC}(\ell,\gamma)\leq\cfrac{C(n,\gamma,p)}{\bigg{(}1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{\frac{n-2}{2}}\bigg{)}^{\frac{2p^{\ast}}{n}\big{(}1-\frac{\gamma}{n}\big{)}}}.$
Theorem 1.6 ( Critical Sharp Singular Adams’ Type Inequality).
Let $1<p<\frac{n}{2}$, $0\leq\gamma<n$, $n\geq 4$ and $a,b>0$. Then there holds
$\text{ATC}_{a,b}(\gamma)=\sup_{\ell\in(0,\beta(n,2))}\Bigg{(}\cfrac{1-\big{(}\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})a}}{(\frac{\ell}{\beta(n,2)}\big{)}^{(\frac{n-2}{n})b}}\Bigg{)}^{\frac{p^{\ast}}{b}\big{(}1-\frac{\gamma}{n}\big{)}}\text{ATSC}(\ell,\gamma).$ (1.15)
To state our main result, we need the following compact embedding, which play crucial rule in the study of the problem ($\boldsymbol{S}$). Its statement can be read in the following way.
Theorem 1.7.
Assume that $1<p<\frac{n}{2}$, $\gamma\in(0,n)$ and $n\geq 4$. Then the embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ is compact for all $\varrho\geq p^{\ast}$, where $p^{\ast}:=\frac{np}{n-2p}$, so-called critical Sobolev exponent.
@@
Suppose that $1<p<\frac{n}{2}$, $\gamma\in(0,n)$ and $n\geq 4$. Let the hypotheses $(g_{1})-(g_{6})$ be satisfied. Moreover, let there exists $\lambda_{0}$ be sufficiently large enough such that $(g_{6})$ holds for $\lambda\geq\lambda_{0}$, then the problem ($\boldsymbol{S}$) has a nontrivial weak solution.
The structure of this article is organized in the following manner: Section 2 is dedicated to some preliminary results, which are needed in our proofs. In Section 3, we shall prove Theorem 1.3 and Theorem 1.4, respectively. Section 4 is devoted to the proof of Theorem 1.5 and Theorem 1.6, whereas in Section 5, we shall provide the proof of Theorem 1.7. Finally, due to the applications of these results, and using the mountain pass theorem, in Section 6, we prove the main result of the article, that is, Theorem 1.8.

### v3

@@ 1. Introduction and main results
for all $m\in\mathbb{N}$ with $m<n$ and $\alpha\in[0,n)$. Consequently, if $\beta>\beta_{\alpha,n,m}$, then the supremum (1.7) is no longer finite. Note that in the second-order Sobolev space $W^{2,\frac{n}{2}}(\mathbb{R}^{n})$ with $n\geq 4$, the sharp singular Adams’ inequality was investigated in [33] under the following constraint
$\bigg\{u\in W^{2,\frac{n}{2}}(\mathbb{R}^{n}):~\|\Delta u\|^{\frac{n}{2}}_{\frac{n}{2}}+\tau\|u\|^{\frac{n}{2}}_{\frac{n}{2}}\leq 1~\text{for all}~\tau>0\bigg\}.$
@@
Later, this result was again generalized in [55, Theorem 1.1] into the arbitrary Sobolev space $W^{m,\frac{n}{m}}(\mathbb{R}^{n})$. Explicitly, they proved a strengthened version of the singular Adams’ inequality as follows: let $\tau>0$, then
@@
(1.8) $\sup_{u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}),~\|\nabla^{m}u\|^{\frac{n}{m}}_{\frac{n}{m}}+\tau\|u\|^{\frac{n}{m}}_{\frac{n}{m}}~\leq 1}~\int_{\mathbb{R}^{n}}\frac{\Phi(\beta|u|^{\frac{n}{n-m}})}{|x|^{\alpha}}\,\mathrm{d}x<+\infty,~\forall~\beta\leq\beta_{\alpha,n,m},$
@@
In recent years, Adams’ type inequalities have been widely studied by many authors across diverse domains such as Hyperbolic spaces, Lorentz spaces, CR spheres, compact Riemannian manifolds, and so on. In this context, we cite some delightful works captured in [28, 27, 26]. For applications of Adams’ inequalities in the analysis of elliptic PDEs, we also refer to some notable articles [24, 15, 8].
It is worth mentioning that the concentration-compactness principle was first studied in higher-order Sobolev spaces in [23, Theorem 1]. Later, on exploiting the Hilbert structure of $W^{2,2}(\mathbb{R}^{4})$, the authors [15] established a sharp singular concentration-compactness principle and, as an application, they studied a class of biharmonic equations in dimension four. Then, it was fully extended to the entire space $\mathbb{R}^{n}$ in [44] and this result can be read as follows: let $u_{k}\rightharpoonup u$ weakly in $W^{m,\frac{n}{m}}(\mathbb{R}^{n})$ with $u\neq 0$, $\|u_{k}\|_{m,\frac{n}{m}}\leq 1$ and
@@
In addition, the author also proved that the above result holds whenever $\|\cdot\|_{m,\frac{n}{m}}$ is replaced by the Ruf norm on $W^{m,\frac{n}{m}}(\mathbb{R}^{n})$, which is defined by
$\|u\|_{\text{Ruf}}:=\bigg(\|u\|^{\frac{n}{m}}_{\frac{n}{m}}+\|\nabla^{m}u\|^{\frac{n}{m}}_{\frac{n}{m}}\bigg)^{\frac{m}{n}},~\forall~u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}).$
Moreover, we also mention that the asymptotic behavior of the supremum for the subcritical sharp Moser-Trudinger inequalities in $\mathbb{R}^{n}$ with different function spaces can be found in recent works [36, 35, 48]. In addition, it was also proved in these articles that the critical and subcritical sharp Moser-Trudinger inequalities are indeed equivalent. In addition, similar results for higher-order Sobolev spaces in $\mathbb{R}^{n}$ can be seen in [34].
Let $n\geq 4$ and $1<p<\frac{n}{2}$ be hold. Next, we define our working function space $E$ as follows
$E:=\Bigg\{u\in L^{1}_{\textit{loc}}(\mathbb{R}^{n})\colon~\int_{\mathbb{R}^{n}}|\Delta u|^{p}\,\mathrm{d}x<+\infty,~\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}}\,\mathrm{d}x<+\infty\Bigg\},$
which is the completion of $C_{0}^{\infty}(\mathbb{R}^{n})$ with respect to the following norm
@@
where $\|\cdot\|_{s}$ standards $L^{s}(\mathbb{R}^{n})$ norm for $s\in[1,+\infty]$. In addition, one can observe that if we define another norm
$\|u\|_{E}:=\|\Delta u\|_{\frac{n}{2}}+\|\Delta u\|_{p},~\forall~u\in E,$
then $\|\cdot\|$ and $\|\cdot\|_{E}$ are equivalent norms for the space $E$. Indeed, we have the following relation
(1.9) $\|u\|\leq\|u\|_{E}\leq 2^{\frac{n-2}{n}}\|u\|,~\forall~u\in E.$
By using standard arguments, we remark that $(E,\|\cdot\|_{E})$ is a uniform convex, reflexive, and separable Banach space. Moreover, since reflexivity is preserved under equivalent norms, we infer that $(E,\|\cdot\|)$ is also a reflexive Banach space. In addition, the space $D^{2,p}(\mathbb{R}^{n})$ is defined as the completion of $C_{0}^{\infty}(\mathbb{R}^{n})$ equipped with the norm
@@
Due to the Sobolev embedding, one can see that the chain of embeddings $E\hookrightarrow D^{2,p}(\mathbb{R}^{n})\hookrightarrow L^{p^{\ast}}(\mathbb{R}^{n})$ holds. Now, for any $\gamma\in[0,n)$, we define the space $L^{r}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$, consisting of all real-valued measurable functions, with $|u|^{r}|x|^{-\gamma}\in L^{1}(\mathbb{R}^{n})$, equipped with the norm
@@
  • $(E^{*},\|\cdot\|_{*})$, $\langle{\cdot,\cdot}\rangle$ denotes the continuous dual of $(E,\|\cdot\|)$,
  • $\langle\cdot,\cdot\rangle$ denotes the duality order pair between $E^{\ast}$ and $E$,
  • $o_{k}(1)$ denotes that the real sequence converges to zero as $k\to\infty$,
@@
Briefly speaking, as our interest is to study a class of $(p,\frac{n}{2})$-biharmonic equations in the space $E$, it should be pointed out that, unfortunately, we cannot use the singular Adams’ inequality that appeared in [33, Theorem 1.5]. Thus, motivated by the above-mentioned works and primarily inspired by the works in [51, 33, 44, 15, 13, 5, 34], we establish a new version of Adams’ inequality in space $E$, its concentration-compactness principle, and as a byproduct of these inequalities, we consider a class of fourth-order elliptic PDEs involving $(p,\frac{n}{2})$-biharmonic operator in the entire space $\mathbb{R}^{n}$. For this, we first define the Young function $\Phi_{\alpha,j_{0}}:\mathbb{R}\to[0,+\infty)$ by
(1.10) $\Phi_{\alpha,j_{0}}(s):=\operatorname{exp}\big(\alpha|s|^{\frac{n}{n-2}}\big)-\sum_{j=0}^{j_{0}-1}\frac{\alpha^{j}|s|^{\frac{nj}{n-2}}}{j!}$
with
$j_{0}:=\left\lceil\frac{p^{\ast}(n-2)}{n}\right\rceil=\min\Bigg\{j\in\mathbb{N}:~j\geq\frac{p^{\ast}(n-2)}{n}\Bigg\},$
where $\left\lceil\cdot\right\rceil$ stands for the ceiling function. Furthermore, we emphasized that our work examines the more generalized singular Adams’ type inequality, which includes the nonsingular situation, that is, $\gamma\in[0,n)$. It can be interpreted as follows.
Theorem 1.1 (Sharp Singular Adams’ Type Inequality).
Let $n\geq 4$, $1<p<\frac{n}{2}$ and $0\leq\gamma<n$ be hold. Then for all $0\leq\alpha\leq\beta_{\gamma,n}:=\big(1-\frac{\gamma}{n}\big)\beta(n,2)$ and $u\in E$, there holds
(1.11) $\sup_{u\in E,~\|u\|\leq 1}\int_{\mathbb{R}^{n}}\frac{\Phi_{\alpha,j_{0}}(u)}{|x|^{\gamma}}\,\mathrm{d}x<+\infty,$
where $\Phi_{\alpha,j_{0}}(\cdot)$ is defined in (1.10) and $\beta(n,2)=n\big[(n-2)\omega^{\frac{2}{n}}_{n-1}\big]^{\frac{n}{n-2}}$ with $\omega_{n-1}$ stands for the measure of the unit sphere in $\mathbb{R}^{n}$. Further, the constant $\beta_{\gamma,n}$ is sharp, that is, if $\alpha>\beta_{\gamma,n}$, then the supremum (1.11) is infinite.
Theorem 1.2 (Sharp Singular Concentration-Compactness Principle).
Let $n\geq 4$, $1<p<\frac{n}{2}$ and $0\leq\gamma<n$ be hold. If $\{u_{k}\}_{k}\subset E$ is a sequence such that $\|u_{k}\|=1$ and $u_{k}\rightharpoonup u\neq 0$ in $E$ as $k\to\infty$, then for all $\ell$ satisfying
we have
(1.12) $\sup_{k\in\mathbb{N}}\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x<+\infty.$
The following natural question is still open at this time:
[…]
Theorem 1.3 (Subcritical Sharp Singular Adams’ Type Inequality).
Let $n\geq 4$, $1<p<\frac{n}{2}$, $0\leq\gamma<n$ and $0\leq\ell<\beta(n,2)$ be hold. Then there holds
The constant $\beta(n,2)$ is sharp in the sense that $\text{ATSC}(\beta(n,2),\gamma)=+\infty$. Also, there exists two positive constants $c(\gamma,p,n)$ and $C(\gamma,p,n)$ such that whenever $l\underset{\approx}{<}\beta(n,2)$, there holds
$\cfrac{c(\gamma,p,n)}{\Bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\Bigg)^{\frac{2p^{\ast}}{n}\big(1-\frac{\gamma}{n}\big)}}\leq\text{ATSC}(\ell,\gamma)\leq\cfrac{C(\gamma,p,n)}{\Bigg(1-\bigg(\frac{\ell}{\beta(n,2)}\bigg)^{\frac{n-2}{2}}\Bigg)^{\frac{2p^{\ast}}{n}\big(1-\frac{\gamma}{n}\big)}}.$
Theorem 1.4 (Critical Sharp Singular Adams’ Type Inequality).
Let $n\geq 4$, $1<p<\frac{n}{2}$, $0\leq\gamma<n$ and $a,b>0$ be hold. Then there holds
(1.13) $\text{ATC}_{a,b}(\gamma)=\sup_{\ell\in(0,\beta(n,2))}\Bigg(\cfrac{1-\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})a}}{\big(\frac{\ell}{\beta(n,2)}\big)^{(\frac{n-2}{n})b}}\Bigg)^{\frac{p^{\ast}}{b}\big(1-\frac{\gamma}{n}\big)}\text{ATSC}(\ell,\gamma).$
Prior to our investigation, we remark that the compact embedding that is explained below is crucial in our analysis. The following is the interpretation of its statement.
Theorem 1.5.
Let $n\geq 4$, $1<p<\frac{n}{2}$ and $\gamma\in(0,n)$ be hold. Then the embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ is compact for all $\varrho\geq p^{\ast}$, where $p^{\ast}:=\frac{np}{n-2p}$, the so-called critical Sobolev exponent.
Moreover, by the applications of Theorem 1.1, Theorem 1.2 and Theorem 1.5, we investigate the existence of nontrivial solutions to a class of $(p,\frac{n}{2})$-biharmonic equations with singular exponential growth in $\mathbb{R}^{n}$. Specifically, we study the following equation
($\mathcal{P}$) $\Delta^{2}_{p}u+\Delta^{2}_{\frac{n}{2}}u=\frac{g(x,u)}{|x|^{\gamma}}\quad\text{in}\quad\mathbb{R}^{n},$
with $n\geq 4$, $1<p<\frac{n}{2}$, $\gamma\in(0,n)$ and $\Delta^{2}_{t}~\cdot:=\Delta(|\Delta\cdot|^{t-2}\Delta~\cdot)$ is a fourth-order operator, which is known as the standard $t$-biharmonic operator for all $t>1$. The nonlinearity $g:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function and has critical exponential growth at infinity, that is, it behaves like $\exp{(\alpha|s|^{\frac{n}{n-2}})}$ as $|s|\to~+\infty$ for some $\alpha>0$, which means that there exists a positive constant $\alpha_{0}$ such that
uniformly with respect to $x\in\mathbb{R}^{n}$. This notion of criticality is basically driven by the well-known Moser-Trudinger type inequality. Throughout this article, without further mention, we assume that $g:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$ satisfies the following properties.
  1. ($\bf g{1}$)
  2. ($\bf g{2}$)
  3. ($\bf g{3}$)
    There exists constant $\alpha_{0}>0$ such that
  4. ($\bf g{4}$)
    There exists constant $\mu>\frac{n}{2}$ such that
  5. ($\bf g{5}$)
    There exists positive constants $s_{0}$ and $M_{0}$ such that for all $x\in\mathbb{R}^{n}$, we have
    $0<G(x,s)\leq M_{0}|g(x,s)|\quad\text{for all}\quad|s|\geq s_{0}.$
  6. ($\bf g{6}$)
In addition, one can notice that, due to ($\bf g{2}$) and ($\bf g{3}$), for any fixed $q>\max\{p^{\ast},\frac{n}{2}\}$ and $\alpha>\alpha_{0}$, there exists $\zeta>0$ and a constant $D_{\zeta}>0$ depending upon $\zeta,~\alpha$ and $q$ such that
(1.14) $|g(x,s)|\leq\zeta|s|^{\tau-1}+D_{\zeta}|s|^{q-1}\Phi_{\alpha,j_{0}}(s)\quad\text{for all}\quad(x,s)\in\mathbb{R}^{n}\times\mathbb{R}$
and by using ($\bf g{4}$), we have
(1.15) $|G(x,s)|\leq\zeta|s|^{\tau}+D_{\zeta}|s|^{q}\Phi_{\alpha,j_{0}}(s)\quad\text{for all}\quad(x,s)\in\mathbb{R}^{n}\times\mathbb{R},$
where $\Phi_{\alpha,j_{0}}(\cdot)$ is the Young function defined in (1.10). Moreover, one can see that the following function satisfies all the above mentioned hypotheses.
Example 1.6.
The function $G\colon\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$ given by $G(x,s):=\lambda|s|^{\vartheta}\operatorname{exp}\big(\alpha_{0}|s|^{\frac{n}{n-2}}\big)$ for all $(x,s)\in\mathbb{R}^{n}\times\mathbb{R},$ where $n\geq 4$, $\lambda>0$, $0<\alpha_{0}<,\alpha$, $\vartheta>\tau,\mu>\max\{p^{\ast},\frac{n}{2}\}$, $g(x,s)=\partial_{s}G(x,s)$ for all $(x,s)\in\mathbb{R}^{n}\times(\mathbb{R}\setminus\{0\})$, and $g(x,0)=0$ for all $x\in\mathbb{R}^{n}$, then $g$ satisfies all the hypotheses ($\bf g{1}$)($\bf g{6}$).
@@
Next, we would like to emphasize the important aspects and novelties of this article below:
@@
    In higher-order Sobolev spaces, the Pólya–Szegö types inequalities don’t hold, which causes us to face several barriers in proving the singular Adams’ type inequality in $E$ and the associated concentration-compactness principle, as can be shown in Theorem 1.1 and Theorem 1.2.
@@
    Due to the critical exponential growth of the nonlinearity, there is a healthy competition between the nonhomogeneous $(p,\frac{n}{2})$-biharmonic operator and the nonlinear term because the embedding $E\hookrightarrow L_{\Phi_{\alpha,j_{0}}}(\mathbb{R}^{n})$ is not compact for the non-singular case, that is, $\gamma=0$, for instance, one can see to Theorem 1.1, where $\Phi_{\alpha,j_{0}}(\cdot)$ is the Young function defined in (1.10). Therefore, proving the compactness of Palais-Smale sequences for the variational energy associated with our main problem becomes a challenge in this situation. Note that a similar type of difficulty also arises for the singular case, that is, $\gamma\neq 0$. To avoid this difficulty, we shall discuss a concentration-compactness principle, that is, Theorem 1.2.
@@
    The study of the existence of solutions to our main problem relies on the compact embedding $E\hookrightarrow L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)$ for all $\varrho\geq p^{\ast}$ and $\gamma\in(0,n)$, thanks to Theorem 1.5. It follows that our method is different from $\gamma=0$. However, we highlighted here that by employing the same method developed in this article and Lemma 2.5, one can study the non-singular case, that is, $\gamma=0$.
@@
    It is well-known that for any $u\in E$, there is no guarantee that $|u|,~u^{\pm}\in E$, where $u^{\pm}=\max\{\pm u,0\}$ and thus, we shall unable to get positive solutions to our main problem.
@@
The following defines the weak formulation of our main problem.
@@
We say that $u\in E$ is a weak solution to the problem ($\mathcal{P}$), if there holds
$\int_{\mathbb{R}^{n}}|\Delta u|^{p-2}\Delta u~\Delta v~\mathrm{d}x+\int_{\mathbb{R}^{n}}|\Delta u|^{\frac{n}{2}-2}\Delta u~\Delta v~\mathrm{d}x=\int_{\mathbb{R}^{n}}\frac{g(x,u)v}{|x|^{\gamma}}\,\mathrm{d}x,~\forall~v\in E.$
@@
Suppose that $1<p<\frac{n}{2}$, $\gamma\in(0,n)$ and $n\geq 4$. Let the hypotheses ($\bf g{1}$)($\bf g{6}$) be satisfied. Moreover, let there exists $\lambda_{0}$ be sufficiently large enough such that ($\bf g{6}$) holds for $\lambda\geq\lambda_{0}$, then the problem ($\mathcal{P}$) has a nontrivial weak solution.
The structure of this article is organized in the following manner: Section 2 is dedicated to some preliminary results, which are needed in our proofs. In Section 3 and Section 4, by employing a rearrangement-free approach and Schwarz symmetrization, we shall prove Theorem 1.1 and Theorem 1.2. Whereas in Section 5 and Section 6, we shall provide the proof of Theorem 1.3 and Theorem 1.4. In addition, we establish a compact embedding result; for instance, see Theorem 1.5 in Section 7. Finally, due to the applications of these results, and using the mountain pass theorem, in Section 8, we prove Theorem 1.8.